Be sure to answer all parts of the question:

a. A fabric remnant from a burial site has a 14C/12C ratio of 0.655 of the original value. How old is the fabric?
____ yr

b. If 5.01 × 10−9 mol of 176W emits 1.07 × 1015 β+ in 1.00 h, what is the decay constant? Enter your answer in scientific notation.
____ × 10___ h−1
Answer #1

(a) The change in the amount of 14C(N) IS given

d(N)/d(t) = -λ*N where λ is decay constant

integration of above equation -

∫dN/N = -λ∫dt ; N = N0e-λt.........(1)

Half life of carbon is 5730 years

it means at t = t1/2 ; N = N0/2 ; Putting these value in equation (1)

N0/2 =N0e-λt  
λ = ln(2) / t1/2 = 0.693/5730=1.21x10-4 yr-1

Given N/N0 = 0.655 Using value of N/N0 and λ in (1)   

N/N0 = e-λt

ln(N0/N) = λt ......(2)

t = ln(1/0.655)/1.21x10-4 yr-1    =3496yr

So the fabric is 3496yrs old

(b) N = N0e-λt

Initially no of particle ( N0)= mole*NA   =5.01 × 10−9*6.023*1023 = 3*1015

Given N(particles(β+) at any time ) = 3*1015-1.07 × 1015= 2*1015

Time(t) = 1hr

By equation (2)

ln(3*1015/2 × 1015) = λ*1hr ; λ = 0.405hr-1

The decay constant( λ ) =4.05*10-1hr-1